### v2

@@
Abstract In this work, we introduce a partitioned optimization framework (POf) to ease the solving process for optimization problems for which fixing some variables to a tunable value removes some difficulties. The variables space may be continuously partitioned into subsets where these variables are fixed to a given value, so that minimizing the objective function restricted to any of the partition sets is easier than minimizing the objective function over the whole variables space. Then our methodology consists in searching for the partition set minimizing the minimum of the objective function restricted to this set. Indeed, a local solution to the original problem is given by computing this partition set and selecting the minimizer of the objective function restricted to this set. We first formalize this framework and study its theoretical guarantees, and we discuss about a derivative-free optimization (DFO) algorithm able to seek for an optimal partition set. Then we focus on a class of composite optimization problems that may be solved using the POf, and we illustrate in details our methodology on four problems from this class. Finally, we highlight the gain in numerical performance provided by the POf on large-dimensional composite problems by comparing it to two popular DFO solvers.
@@ 1 Introduction
$\underset{\begin{array}[]{c}y\in\mathbb{Y}\end{array}}{\operatorname*{minimize}}\quad\varphi(y)\quad\mbox{subject to}\quad y\in\Omega,$ ($\mathbf{P}_{\texttt{ini}}$)
where $\mathbb{Y}$ is a (possibly infinite-dimensional) normed space, and $\varphi:\mathbb{Y}\to\overline{\mathbb{R}}$ is a (possibly discontinuous) function with nonempty effective domain $Y\triangleq\{y\in\mathbb{Y}:\varphi(y)\neq\pm\infty\}$, and $\Omega\subseteq\mathbb{Y}$ intersects $Y$. We introduce a partitioned optimization framework (POf) to solve Problem ($\mathbf{P}_{\texttt{ini}}$) in cases where $\mathbb{Y}$ may be partitioned in a way that the restriction of the problem to any partition set is significantly simpler to solve than the problem itself. Section 1.1 illustrates the POf, Section 1.2 discusses our motivation to develop it, and Section 1.3 summarizes our contribution and outlines this work.
Notation: We denote by $\overline{\mathbb{R}}\triangleq\mathbb{R}\cup\{\pm\infty\}$$\overline{\mathbb{N}}\triangleq\mathbb{N}\cup\{+\infty\}$$\mathbb{R}^{*}\triangleq\mathbb{R}\setminus\{0\}$ and ¯¯$\overline{\mathbb{N}}^{*}\triangleq\overline{\mathbb{N}}\setminus\{0\}$. Consider two normed spaces $\mathbb{A}$ and $\mathbb{B}$. A set $A\subseteq\mathbb{A}$ is said to be ample [2] if $A\subseteq\mathrm{cl}(\mathrm{int}(A))$ (where $\mathrm{cl}$ and $\mathrm{int}$ denote respectively the closure and interior operators), and to be a continuity set of a function $f:\mathbb{A}\to\overline{\mathbb{R}}$ if $f(a)$ if finite for all $a\in A$ and $f_{|A}$ is continuous. For all collections $(A_{i})_{i=1}^{N}$ of subsets of $\mathbb{A}$$N\in\overline{\mathbb{N}}^{*}$, their union is denoted by $\sqcup_{i=1}^{N}A_{i}$ when the sets are pairwise disjoint. For all sequences $(b^{k})_{k\in\mathbb{N}}\in\mathbb{B}^{\mathbb{N}}$, we denote by $\mathrm{acc}((b^{k})_{k\in\mathbb{N}})$ the set of all accumulation points of $(b^{k})_{k\in\mathbb{N}}$. For all functions $g:\mathbb{A}\to\mathbb{B}$ and all points $a\in\mathbb{A}$, we denote by $\mathcal{ACC}(g;a)$ the union of $\mathrm{acc}((g(a^{k}))_{k\in\mathbb{N}})$ over all sequences $(a^{k})_{k\in\mathbb{N}}\in\mathbb{A}^{\mathbb{N}}$ converging to $a$. For all functions $f:\mathbb{A}\to\overline{\mathbb{R}}$ and all sets $A\subseteq\mathbb{A}$, we denote by $\overline{f_{A}}:\mathbb{A}\to\overline{\mathbb{R}}$ the extreme barrier function [3] mapping all $a\in\mathbb{A}$ to $f(a)$ if $a\in A$ and to +$\infty$+ ∞ if $a\notin A$. In particular, $\overline{\varphi_{\Omega}}$ denotes the extreme barrier function associated to $\varphi$ and $\Omega$ from Problem ($\mathbf{P}_{\texttt{ini}}$).
@@ 1.1 Illustrative example with a composite greybox problem
This section illustrates the POf with no constraints ($\Omega\triangleq\mathbb{Y}$), two variables ($\mathbb{Y}\triangleq\mathbb{R}^{2}$), and
@@
shown in Figure 1 (this example is formally solved in Section 4.1). Note that $\varphi$ is explicitly defined with respect to the functions $\sigma$ and $\varepsilon$, while $\sigma$ and $\varepsilon$ are blackboxes since their expressions are not provided. A possible application of the POf is to exploit partial information in similar contexts.
Figure 1: The function $\varphi$ from Section 1.2, minimized efficiently via the POf. For readability, at each $y_{1}\in[-4,4]$ the two graphs actually plot only the values of the quadratic $y_{2}\mapsto\varphi(y_{1},y_{2})$ for $y_{2}\in[\sigma(y_{1})-1.3,\ \sigma(y_{1})+1.3]$.
Minimizing $\varphi$ over $\mathbb{Y}$ is difficult, remark that $\varphi$ has a simple expression when $y_{1}$ is fixed: for all $y_{1}\in\mathbb{R}$, the function $y_{2}\mapsto\varphi(y_{1},y_{2})$ is quadratic. The first step of the POf exploits this remark by partitioning $\mathbb{Y}$ as $\mathbb{Y}\triangleq\sqcup_{x\in\mathbb{X}}\mathbb{Y}(x)$, where $\mathbb{Y}(x)\triangleq\{x\}\times\mathbb{R}$ for all $x\in\mathbb{X}\triangleq\mathbb{R}$, so that minimizing $\varphi$ over $\mathbb{Y}(x)$ would be simple: in this case, we directly observe that the solution is $(x,\sigma(x))$, with objective function value $\varepsilon(x)$. Then, the POf defines an oracle function $\gamma:\mathbb{X}\to\mathbb{Y}$ that, given any partition index $x\in\mathbb{X}$, computes the minimizer of $\varphi$ over $\mathbb{Y}(x)$; in this case $\gamma(x)\triangleq(x,\sigma(x))$ is analytically tractable for all $x\in\mathbb{X}$. The second step of the POf calculates an optimal partition set index $x\in\mathbb{X}$ minimizing the minimizer of $\varphi$ over the set $\mathbb{Y}(x)$, that is, minimizing $\Phi(x)\triangleq\varphi(\gamma(x))$ (and, in this case, $\Phi(x)=\varepsilon(x)$). This problem is solved by an usual blackbox optimization approach, where we compute $\gamma(x)$ on-the-fly during the optimization process at points $x\in\mathbb{X}$ we seek to evaluate. Finally, the last step of the POf recovers a solution to Problem ($\mathbf{P}_{\texttt{ini}}$), given by $\gamma(x^{*})$, from an optimal partition set $x^{*}$ (in this example, $x^{*}\triangleq 0$ and $\gamma(x^{*})\triangleq(0,0)$). For completeness, define also the index function $\chi$, defined for all $y\triangleq(y_{1},y_{2})\in\mathbb{Y}$ as the unique index $x\in\mathbb{X}$ of the unique partition set $\mathbb{Y}(x)$ containing $y$. In this context, $\chi(y)\triangleq y_{1}$ for all $y\in\mathbb{Y}$, since $y\in\mathbb{Y}(x)$ only when $x=y_{1}$.
@@ 1.2 Motivation for this work
This work originates from the corresponding author’s PhD thesis [9, Chapters 5 and 6] (in French). The initial motivation was to solve an optimal control problem [9, Chapter 7] with discontinuities, which justifies $\mathbb{Y}$ being possibly infinite-dimensional and $\varphi$ being possibly discontinuous. An early version of [9, Chapter 7], without the theory supporting the POf, appears in a technical report [10]. In all cases considered in [10]$\mathbb{Y}$ is an infinite-dimensional functional space and $\varphi$ is discontinuous, while the POf identifies a partition of $\mathbb{Y}$ such that $\dim(\mathbb{X})\leq 4$ and the problem to minimize $\varphi_{|\mathbb{Y}(x)}$ for any $x\in\mathbb{X}$ is an usual optimal control problem that may be solved by most dedicated techniques.
The POf allows to solve several problems that share similarities with the example from Section 1.1. What matters is that a well-chosen partition of $\mathbb{Y}$ allows for a significant simplification of the problem, so that a global solution to the problem restricted to any partition set is tractable (analytically of through a dedicated optimization algorithm) whenever one needs to calculate it. Then the problem is solved by seeking for the optimal partition set index, which is done as an usual blackbox problem (likely low-dimensional even when $\mathbb{Y}$ is infinite-dimensional). A possible field list of applications reads as composite greybox problem where the inputs of the blackbox components of the problem depend only on a few nonlinear combinations of all variables, as discussed in details in Section 4. Others possible applications are completely explicit problems (admitting no blackbox components) that consist in two distinct sub-problems that are connected through a few variables, or are discontinuous but the discontinuities are affected by a few variables only, or become smooth when a constraint linking the variables is added. In particular, our initial motivational problem [10] matches all these examples.
@@ 1.3 Contribution and outline of this work
Our contribution is threefold. First, we formalize the POf and we prove that an optimal partition set index allows to recover a solution to Problem ($\mathbf{P}_{\texttt{ini}}$) under a minimal assumption. Second, we provide an algorithm that computes an optimal partition set under an additional assumption. Third, we empirically measure the performance gain provided by the POf through numerical experiment, by comparing this approach to two state-of-the-art DFO solvers solving Problem ($\mathbf{P}_{\texttt{ini}}$) directly.
Let us stress that ensuring the accessibility of the oracle function $\gamma$ is not within our scope. For ease of presentation and theoretical analysis, we design the POf as if we possess an idealized optimization method that may, for all partition set indices $x$, compute a global minimizer of the restriction $\varphi_{|\mathbb{Y}(x)\cap\Omega}$. Then $\gamma$ consists in the function that, given any $x$, calls this method and returns its output. Section 6 discusses about this restrictive setting and how we plan to significantly relax it in a future work. Similarly, how to find a relevant partition of $\mathbb{Y}$ in general is not within our scope either, since this is problem-dependent. Nevertheless, Section 4 shows how to find such a partition in a class of composite problems, and [9, Chapter 7] and [10] provide a similar guideline in a class of optimal control problems.
@@
This work is organized as follows. A literature review is provided in Section 2. Section 3 addresses our first two goals: it formalizes the POf and highlights its properties in Theorem 1, and it formalizes our strategy to solve Problem ($\mathbf{P}_{\texttt{ini}}$) using the POf in Theorem 2. Our third goal is considered in Sections 4 and 5. Section 4 highlights how Theorem 2 allows to solve some composite optimization problems, and shows that numerical observations match our theoretical claims. Section 5 alters the problems from Section 4 into more challenging versions with a large dimension, and shows that solving these problems with two DFO solvers from the literature (solving the non-reformulated problems) is less efficient than our strategy. Section 6 discusses about the POf and its possible extensions.
2 Literature review
The POf is not suited for blackbox optimization problems but is related to greybox optimization [1]. Indeed, choosing a relevant partition for $\mathbb{Y}$ requires information on the problem, and for all $x\in\mathbb{X}$, ensuring the tractability of $\gamma(x)$ requires that the restriction $\varphi_{|\mathbb{Y}(x)\cap\Omega}$ is simple enough to minimize. Nevertheless, most DFO algorithms [22] tackling greybox problems are initially designed for blackbox contexts, because fine-tuning strategies to greybox cases are problem-dependent. Then, we compute an optimal partition set using an algorithm with a covering step [2] since to our best knowledge only these algorithms are guaranteed to return a local solution in discontinuous contexts (under a mild assumption). We provide ideas in Remark 2 to specialize an algorithm to the context of the POf.
Solving Problem ($\mathbf{P}_{\texttt{ini}}$) directly with an algorithm from derivative-free optimization (DFO) is likely inefficient, since the dimension of $\mathbb{Y}$ may be high while DFO algorithms usually handle up to a few dozens of variables [5, Section 1.4]. For example, both the NOMAD [6] and PRIMA [30] solvers state a critical value of $50$ variables. Some dimension-selection techniques [17] handle large-scale problems with active subspaces [14, 29], or via random directions [12, 19, 26]. These techniques likely outperform the POf when the objective function has a small sensitivity with respect to most of the variables, but likely fail on cases where the objective function has a small sensitivity with respect to each variables individually but has a high sensitivity with respect to all variables jointly. For example, this happens in some optimal control problems [9, Chapter 7]: an alteration of any single state variable has a small impact, but an alteration of all state variables together significantly alters the objective value.
@@ 2 Literature review
The POf is related to parametric optimization [7, 28], which studies problems where the objective function and the feasible set depend on a tunable parameter. In our case, the parametric problem would be the minimization of $\varphi$ over $\mathbb{Y}(x)\cap\Omega$, where the parameter of the problem would be the partition set index $x\in\mathbb{X}$. Any problem fitting in any of the two frameworks may fit in the other. The local continuity of $\gamma$ in a discontinuous context is studied in [18], but, to the best of our knowledge, no reference in parametric optimization claims results similar to ours. Moreover, the terminology partitioned optimization highlights that Problem ($\mathbf{P}_{\texttt{ini}}$) is non-parametric and that the nature of the so-called parameter in the restricted problem (the index of the associated partition set of $\mathbb{Y}$, which results from the chosen partition of $\mathbb{Y}$) is left to the user.
The POf also shares similarities with bilevel optimization [15], which describes problems in which some of the variables are fixed as a solution to a nested subproblem. Any problem expressed in any of the frameworks may be transferred to the other. However, the difference in terminology highlights a contextual difference since Problem ($\mathbf{P}_{\texttt{ini}}$) has only one level and, unlike in usual bilevel problems, there is no conflictual min-max structure. Some references study bilevel nonsmooth problems [25] and apply DFO algorithms [16] but, to the best of our knowledge, none provides results similar to ours.
The terminology partitioned optimization also describes the DFO methods [20, 24]. However, these methods consider a discrete partition which do not simplifies the restricted problem, as they partition an hyperrectangle into smaller ones and recursively partition the most promising. Then, we believe that our continuous and structure-simplifying partition fits in the same terminology without confusion.
@@
The POf may also be related to optimization on manifold [11], which considers problems with a constraint inducing an explicit manifold on the variables space. This happens in our case to the problem restricted to any partition set, when each partition set of $\mathbb{Y}$ is a manifold. In practice, when $\mathbb{Y}$ has a finite dimension, it is likely that $\mathbb{Y}$ is indeed partitioned into a continuum of manifolds.

### v3

@@
Abstract This work tackles a class of optimization problems where fixing some well-chosen combinations of the variables makes the problem substantially easier to solve. We consider that the variables space may be partitioned into subsets that fix these combinations to given values, so the restriction of the problem to any of the partition sets admits a tractable solution. Then, we exhibit a reformulation of the original problem that consists in searching for the partition set index that minimizes the objective value of the solution to the restricted problem. We name partitioned optimization framework (POf) the formalization of this class of problems and this reformulated problem. As we prove in this work, the POf allows solving the original problem by focusing on the reformulated problem: all solutions to the reformulated problem are partition indices for which the solution to the associated restricted problem is also a solution to the original problem. Second, we introduce a derivative-free partitioned optimization method (DFPOm) to efficiently solve problems that fit in the POf. We prove that the reformulated problem is nicely handled by a class of derivative-free optimization (DFO) algorithms named algorithms with a covering step. Then the DFPOm consists in solving the reformulated problem using such DFO algorithm with a covering step to obtain an optimal partition index, and to return the solution to the associated restricted problem as a solution to the initial problem. Finally, we illustrate the DFPOm on a class of problems called composite greybox problems, and we highlight the gain in numerical performance provided by the DFPOm on large-dimensional instances by comparing it to two popular DFO solvers.
@@ 1 Introduction
$\underset{\begin{array}[]{c}y\in\mathbb{Y}\end{array}}{\operatorname*{minimize}}\quad\varphi(y)\quad\mbox{subject to}\quad y\in\Omega,$ ($\mathbf{P}_{\texttt{i}}$)
where $\mathbb{Y}$ is a (possibly infinite-dimensional) Banach space, and $\varphi:\mathbb{Y}\to\overline{\mathbb{R}}$ (possibly discontinuous) has a nonempty effective domain $Y\triangleq\{y\in\mathbb{Y}:\varphi(y)\neq\pm\infty\}$, and $\Omega\subseteq\mathbb{Y}$ (possibly not closed) intersects $Y$. We consider that $\mathbb{Y}$ may be partitioned so that restricting Problem ($\mathbf{P}_{\texttt{i}}$) to any partition set makes it significantly simpler to solve. We introduce a partitioned optimization framework (POf) to study this context, and derivative-free partition optimization method (DFPOm) to efficiently solve Problem ($\mathbf{P}_{\texttt{i}}$) when it fits in the POf. Section 1.1 illustrates the POf, Section 1.2 discusses our motivation to develop it, and Section 1.3 summarizes our contribution and outlines this work.
@@ 1.1 Illustrative example with a composite greybox problem
This section illustrates the POf with two variables ($\mathbb{Y}\triangleq\mathbb{R}^{2}$), no constraints ($\Omega\triangleq\mathbb{Y}$), and
@@
shown in Figure 1 (this example is solved in Section 4.1). Note that $\varphi$ is explicitly defined with respect to the functions $\sigma$ and $\varepsilon$, while $\sigma$ and $\varepsilon$ are blackboxes [6] since their expressions are not provided.
Figure 1: The function $\varphi$ from Sections 1.2 and 4.1, minimized efficiently via the POf. For readability, at each $y_{1}\in[-4,4]$ the two graphs actually plot only the values of the quadratic $y_{2}\mapsto\varphi(y_{1},y_{2})$ for $y_{2}\in[\sigma(y_{1})-1.3,\ \sigma(y_{1})+1.3]$.
Minimizing $\varphi$ over $\mathbb{Y}$ is difficult. Yet, $\varphi$ has a simpler expression when $y_{1}$ is fixed: for all $y_{1}\in\mathbb{R}$, the function $y_{2}\mapsto\varphi(y_{1},y_{2})$ is quadratic. We exploit this property by partitioning $\mathbb{Y}$ as $\mathbb{Y}=\sqcup_{x\in\mathbb{X}}\mathbb{Y}(x)$, where for all partition set index $x\in\mathbb{X}\triangleq\mathbb{R}$, the partition set $\mathbb{Y}(x)\triangleq\{x\}\times\mathbb{R}$ is such that the restriction of $\varphi$ to $\mathbb{Y}(x)$ (denoted by $\varphi_{|\mathbb{Y}(x)}$) is simple to minimize. We define an oracle function $\gamma:\mathbb{X}\to\mathbb{Y}$ that, for all $x\in\mathbb{X}$, computes the minimizer of $\varphi_{|\mathbb{Y}(x)}$. Its associated objective function value is denoted by $\Phi(x)\triangleq\varphi(\gamma(x))$. As we highlight in this work, the optimal partition set index $x^{*}$ minimizing $\Phi$ provides a solution to the initial problem as the point $\gamma(x^{*})$. In this example, for all $x\in\mathbb{X}$, the minimizer of $\varphi_{|\mathbb{Y}(x)}$ is $\gamma(x)\triangleq(x,\sigma(x))$, and its objective function value equals $\Phi(x)=\varepsilon(x)$. The minimizer $x^{*}\triangleq 0$ of $\Phi$ yields $\gamma(x^{*})\triangleq(0,0)$, which is indeed the global solution to the problem.
The POf formalizes the class of optimization problems where a similar approach is possible. We also introduce our DFPOm to compute the optimal partition set $x^{*}$, since the problem of minimizing $\Phi$ is not trivial in general. For completeness, let us also introduce another function the POf relies on, the index function $\chi$, defined for all $y\in\mathbb{Y}$ as the unique index $x\in\mathbb{X}$ of the unique partition set $\mathbb{Y}(x)$ containing $y$. In the example above, $\chi(y)\triangleq y_{1}$ for all $y\triangleq(y_{1},y_{2})\in\mathbb{Y}$, since $y\in\mathbb{Y}(x)$ only when $x=y_{1}$.
@@ 1.2 Motivation for this work
This work introduces the POf as a new optimization framework that leverages inherent structure of some optimization problems to describe their potential solutions, and the DFPOm to exploit the POf in practice. Several applications may enjoy the DFPOm.
A first broad field of applications is optimal control problems [32] with discontinuous Mayer cost. An example consists in optimizing the flight of a parachute so that it lands on a target made of concentric rings with different costs, while minimizing overall efforts. The Mayer cost of the flight is the cost associated to the ring the parachute ends in, which is therefore discontinuous. However, if we impose the landing point of the parachute, the Mayer cost becomes fixed, so the problem becomes usual, and then solvable (at least numerically with standard methods). Thus, we may partition the variables space (the full trajectory of the parachute plus the control to ensure it) according to the landing point. This context was our initial motivation in developing the POf and to consider $\mathbb{Y}$ infinite-dimensional. This study is detailed in our technical report [11] and in the PhD thesis [10, Chapter 7] (in French).
Another rich field is composite greybox problem [21, 23] where, as in Section 1.1, the inputs of the blackbox components of the problem are low-dimensional. Such problems are discussed in Section 4.
Finally, we have initiated applications to counterfactuals [31, 33] for contextual optimization [28]. Computing such counterfactuals is a trending topic in machine learning, and its relation to the POf is natural. A contextual problem consists in minimizing an objective function parameterized by a context vector, that is usually predicted from data by some machine learning model. Given a context $x$, its associated contextual solution $y^{*}(x)$, and another element $y\neq y^{*}(x)$, a counterfactual of $x$ with respect to $y$ is a context $x^{\prime}\neq x$ close to $x$ and such that $y^{*}(x^{\prime})=y$.
Others possible applications may likely connect to the POf, such as almost separable problems [29] and problems that become smooth when a constraint linking the variables is added.
@@ 1.3 Contribution and outline of this work
Our contribution is threefold. First, we state the POf, and we show that we may recover a solution to Problem ($\mathbf{P}_{\texttt{i}}$) from an optimal partition set index. Second, we propose the DFPOm to exploit the POf, using a class of derivative-free optimization (DFO) algorithms to compute an optimal partition set index. Third, we empirically assert the performance of the DFPOm through numerical experiments comparing it against two state-of-the-art DFO solvers solving Problem ($\mathbf{P}_{\texttt{i}}$) directly.
@@
This work is organized as follows. A literature review is provided in Section 2. Section 3 addresses our first two goals: it formalizes the POf and highlights its properties in Theorem 1, and it derives the DFPOm to solve Problem ($\mathbf{P}_{\texttt{i}}$) in Theorem 2. Our third goal follows in Sections 4 and 5. Section 4 shows how the DFPOm solves some composite optimization problems, and that numerical observations match our theoretical claims. Section 5 considers more challenging problem instances, and shows that solving these problems directly with two DFO solvers from the literature is less efficient than our strategy. Finally, Section 6 discusses our work and its possible extensions.
2 Review of related literature
@@
The POf relates to the frameworks of parametric optimization [8, 30] and bilevel optimization [15]. Indeed, parametric optimization studies problems where the objective function and the feasible set depend on a tuneable parameter. In our case, the parameterized problem is the minimization of $\varphi$ over $\mathbb{Y}(x)\cap\Omega$, where the parameter of the problem is the partition set index $x\in\mathbb{X}$. The local continuity of the parametric solution in discontinuous contexts is studied in [18], but, to the best of our knowledge, no reference in parametric optimization claims results similar to ours. Moreover, the terminology partitioned optimization highlights that Problem ($\mathbf{P}_{\texttt{i}}$) is non-parametric and that the nature of the so-called parameter in the restricted problem is left to the user (as the indices of the partition set result from the chosen partition of $\mathbb{Y}$). Similarly, bilevel optimization describes problems in which some of the variables are fixed as a solution to a nested subproblem. However, the difference in terminology highlights a contextual difference since Problem ($\mathbf{P}_{\texttt{i}}$) has only one level and, unlike in usual bilevel problems, there is no conflictual min-max structure. Some references study bilevel nonsmooth problems [25] and apply DFO algorithms [16] but, to the best of our knowledge, none provides results similar to ours.
An important aspect of the POf is its exploitation through the DFPOm. The problem of computing an optimal partition set index is akin to a usual blackbox optimization [1] problem. To solve it, we use DFO algorithms with a covering step [3], since to our best knowledge they are the only algorithms guaranteed to return a local solution in discontinuous contexts under mild assumptions.
Although the DFPOm leverages DFO, we remark that the POf itself targets problems that differ from usual use-cases from DFO. Indeed, the dimension of $\mathbb{Y}$ may be high in Problem ($\mathbf{P}_{\texttt{i}}$), while DFO algorithms usually handle up to a few dozens of variables [6, Section 1.4]. Moreover, the POf requires some information on the problem to define a relevant partition of $\mathbb{Y}$ and ensure the tractability of the minimizer of $\varphi$ over each partition set, which is not available in the blackbox optimization setting that is popular in DFO. The DFO contexts that match the most with the POf are greybox optimization [2] and large-dimension concerns. Greybox optimization considers problems where some information on the problem is available, as in our example in Section 1.2. Regarding large-dimensional DFO, the POf has complementarity with the literature. Indeed, usual techniques rely on dimension-selection [17] that assume active subspaces [14, 34], or random directions [13, 19, 26]. These techniques likely outperform the POf when the objective function has a small sensitivity with respect to most of the variables but a high sensitivity with respect to the few others. Yet, they likely fail on cases where the objective function mostly depends on a few complex combinations of the variables instead of all variables directly. In contrast, this last context is favourable to the POf. As illustrated in Section 5, the DFPOm outperforms usual DFO algorithms on four problems with this property.
As detailed in Section 4, the main application of the POf that we consider in this work relates to greybox composite optimization problems. Although the associated literature is rich, [21, 23] claim that most papers assume specific structures for a composite function $\varphi\triangleq\widetilde{\varphi}\circ\sigma+\varepsilon$. We observe that our approach has complementarity with those in [21, 23]. Indeed, [21] considers cases where $\widetilde{\varphi}$ is cheap to evaluate, but with a possibly complex structure and few inputs, and $\sigma$ and $\varepsilon$ expensive with possibly many inputs (around $65$ at most). In contrast, we better tackle cases where $\widetilde{\varphi}$ has a simple structure and numerous inputs but $\sigma$ and $\varepsilon$ have few inputs. Also, [23] also considers $\widetilde{\varphi}$ well-structured and cheap to evaluate, but unlike us, it assumes that $\sigma$ and $\varepsilon$ are smooth.
@@
The terminology partitioned optimization also describes the DFO methods [20, 24]. However, these methods consider a discrete partition which do not simplify the restricted problem, as they partition a hyperrectangle into smaller ones and recursively partition the most promising. Then, we believe that our continuous and structure-simplifying partition fits in the same terminology without confusion. The POf may also be related to optimization on manifold [12], which considers problems with a constraint inducing an explicit manifold on the variables space. This happens in our case to the problem restricted to any partition set, when each partition set of $\mathbb{Y}$ is a manifold. In practice, when $\mathbb{Y}$ has a finite dimension, it is likely that $\mathbb{Y}$ is indeed partitioned into a continuum of manifolds.